OpenStudy (anonymous):

sqrt(x^2 - 9) = x - 1

OpenStudy (anonymous):

\[\sqrt{x ^{2}-9}=x-1\]

OpenStudy (anonymous):

square both sides to get x^2-9=(x-1)^2

OpenStudy (anonymous):

factor out (x-1)^2 to get x^2-2x+1

OpenStudy (anonymous):

*facepalm* I had that but thanks I just forgot to subtract x^2

OpenStudy (anonymous):

lol no prod :D

OpenStudy (anonymous):

prob* lol